A vector has a magnitude of 9.1 and is in quadrant 1 with an angle of 30.3° with respect to the +y-axis. What is the ×-component of the vector?


1
Expert's answer
2021-03-25T19:28:47-0400

We can find the "x"-component of the vector as follows:


"sin\theta=\dfrac{A_x}{A},""A_x=Asin\theta=9.1\cdot sin30.3^{\circ}=4.6."